OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______ Numerical Answers Expected! Pleeease help me! :)

OpenStudy (anonymous):

f(3) = f(1) + f(2) + f(2) = 2 + 5 + 5 = 12 f(4) = f(1) + f(2) + f(3) = 2 + 5 + 12 = 19 f(5) = f(1) + f(2) + f(4) = 2 + 5 + 19 = 26 <--

OpenStudy (anonymous):

go to www.wolframalpha.com for your equation help. they will give u steps along with answer

OpenStudy (anonymous):